# LANTERN: TCR-Peptide Binding Prediction via Large Language Model Representations

### v1

@@ 4.1 Benchmark Evaluation of LANTERN
The objective is to predict whether the TCR and peptide bind with each other. In LANTERN, we choose Molformer as the peptide SMILES pretraining model and use other state-of-the-art models as baselines for performance comparison. These include Large Language Model-based approaches, including Smiles-Bert [36], TEINet [14], and ChemBERTa [4], as well as end-to-end deep learning models such as TITAN [37], ERGO II [33], NetTCR [26], Dlptcr [42], and Imrex [27]. The main results and analysis demonstrate the powerful performance of Large Language Models in zero-shot learning, effectively capturing the interaction between TCR and peptide. Additionally, their few-shot learning capabilities highlight their ability to remember and generalize knowledge from other domains to the binding prediction problem.
4.2 Performance of LANTERN Compared to Baseline Models
@@
Figure 2: Performance of LANTERN on four datasets. | a. AUC-ROC scores across models for NA dataset. b. AUC-ROC scores across models for RN dataset. c. Accuracy scores across models for the four datasets.
@@
LANTERN exhibits significant improvement over traditional methods. Our experiments highlighted the performance of the LANTERN framework, particularly in the context of the random control setting. Figure 2 presents the AUC-ROC score across the four settings, for the detailed results of all metrics, please refer to the Supplementary Data Table 3 to Table 6. We use six metrics to evaluate performance: AUC-ROC, Accuracy, Recall, Precision, F1-Score, and AUC-PR. Among these, we primarily focus on AUC-ROC, as it is robust to threshold values in binary classification tasks.
@@
4.2.1 Comparison of Generated vs. Original Datasets
Comparing results between Gen NA (Supplementary Data Table 3) and Original NA (Supplementary Data Table 5), and Gen RN (Supplementary Data Table 4) and Original RN (Supplementary Data Table 6) reveal that models trained on generated datasets generally performed better than those trained on original datasets. For instance, LANTERN achieves an AUC-ROC score of 0.88 on Gen NA versus 0.901 on Original NA, and 0.66 on Gen RN versus 0.57 on Original RN. This suggests that the novel negative sample generation method improved model training by providing more realistic and challenging scenarios, closely mimicking the complexities of real-world data. The generation of negative samples in Gen NA and Gen RN datasets allowed for a more balanced and comprehensive training process. This method ensures that the model encounters a variety of binding and non-binding interactions, enhancing its ability to distinguish between them. By incorporating a diverse range of negative samples, the model becomes more adept at recognizing subtle differences in binding patterns, leading to improved prediction accuracy and robustness.
4.2.2 Generally Challengeable Random Control Setting
@@
4.2.3 Embedding Cluster of Pretraining Model
Figure 3: t-SNE visualization of the embeddings from LANTERN model. | a. With pretrain of 0.857 h score. b. Without pretrain of 0.607 h score. c. Show frequency of the pretrained embeddings. d. Show frequency of the no pretrained embeddings.
@@ 4.2.3 Embedding Cluster of Pretraining Model
4.2.4 Quantitative Evaluation of Embedding Performance
4.2.5 Embedding Cluster of Pretraining Model
@@
To further quantify the performance of the embeddings learned by the pretraining model, we calculated several clustering performance metrics: Silhouette Index (SI) and Calinski-Harabasz Index (CHI). Additionally, we evaluated unsupervised metrics, including the Hopkins Score (HS) and Cluster Tendency Score (CTS) [43]. For SI and CH, we employed the K-means clustering method on the embeddings with k = 2. We put the algorithm for these metrics in the supplementary section 2. The results are illustrated in the Figure 4. In Figure 4, we observe the following metrics:
@@
4.3 Few-Shot Learning Capabilities of LANTERN
Figure 5: Few-Shot Learning performance of LANTERN | a. Performance comparison of models over different numbers of seen pairs. b. Performance comparison of models over different seen ratios.
@@ 4.3 Few-Shot Learning Capabilities of LANTERN
5 Discussion and Conclusion
The LANTERN model presents a significant advancement in the field of T-cell receptor (TCR) and peptide-major histocompatibility complex (pMHC) binding prediction, leveraging large language model (LLM) representations to address limitations in current methodologies. The integration of pre-trained embeddings from protein language models and SMILES encoding of peptides provides a nuanced understanding of TCR-pMHC interactions, which has been demonstrated through the model’s ability to perform well in both zero-shot and few-shot learning scenarios. This capability is particularly important in the context of predicting previously unseen epitopes, a key challenge for existing models.
One of the critical strengths of LANTERN is its ability to generalize from minimal data, effectively tackling the issue of data scarcity that plagues many deep learning approaches in immunogenetics. The performance of LANTERN across different negative sampling strategies, particularly in the random control setting, illustrates its robustness and versatility. The model consistently outperformed baseline methods, including traditional deep learning and other LLM-based models, across multiple evaluation metrics such as AUC-ROC, precision, and recall. This indicates that the LANTERN framework captures more complex binding patterns compared to other methods, providing a more accurate and reliable prediction of TCR-pMHC interactions.
However, there is still room for improvement, especially in the random control setting, where the performance, while strong, did not reach the same levels as in the reference control setting. This suggests that further refinement of the negative sample generation process, perhaps incorporating domain-specific knowledge or more sophisticated biological features, could enhance the model’s ability to distinguish between true negative and binding pairs. Future work could explore the use of ensemble learning techniques or the incorporation of multimodal data, such as structural or contextual biological information, to improve the model’s performance further.
@@ 5 Discussion and Conclusion
{credits}
5.0.1 Acknowledgements
We would like to express our gratitude to our collaborators for their insightful discussions and technical support throughout this work.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.

### v2

@@ Results
The objective is to predict whether the TCR and peptide bind with each other. In LANTERN, we choose MolFormer as the peptide SMILES pretraining model and use other state-of-the-art models as baselines for performance comparison. These include Large Language Model-based approaches, including Smiles-Bert ref39, TEINet ref18, and ChemBERTa ref40, as well as end-to-end deep learning models such as TITAN ref15, ERGO II ref7, NetTCR ref25, Dlptcr ref41, and Imrex ref42. The main results and analysis demonstrate the powerful performance of Large Language Models in zero-shot learning, effectively capturing the interaction between TCR and peptide. Additionally, their few-shot learning capabilities highlight their ability to remember and generalize knowledge from other domains to the binding prediction problem.
The model is trained end-to-end using the Adam optimizer with early stopping based on validation AUC. All pretrained encoders (ESM and MolFormer) are frozen or fine-tuned depending on experimental settings. The MHCA module and MLP layers are trained from scratch. The model’s performance is rigorously validated against a held-out test set, employing metrics such as AUC-ROC to quantify prediction accuracy and reliability. This methodology of LANTERN, integrating the powerful ESM-1b and MolFormer models, sets the stage for profound advancements in understanding and predicting the dynamics of protein-ligand interactions, with significant implications for drug discovery and biochemical research.
Performance of LANTERN Model
@@
Figure 2: Performance of LANTERN on four datasets. — a. AUC-ROC scores across models for NA dataset. b. AUC-ROC scores across models for RN dataset. c. Accuracy scores across models for the four datasets.
@@
LANTERN exhibits significant improvement over traditional methods. Our experiments highlighted the performance of the LANTERN framework, particularly in the context of the random control setting. Figure 2 presents the AUC-ROC score across the four settings, for the detailed results of all metrics, please refer to the Table 3 to Table 6. We use six metrics to evaluate performance: AUC-ROC, Accuracy, Recall, Precision, F1-Score, and AUC-PR. Among these, we primarily focus on AUC-ROC, as it is robust to threshold values in binary classification tasks.
Table 3. Result of LANTERN and baseline models on Gen NA
Model AUC-ROC ACC RECALL PRECISION F1-SCORE AUC-PR
LANTERN 0.882 0.728 0.341 0.992 0.483 0.991
Smiles-Bert 0.659 0.702 0.621 0.843 0.811 0.808
TEINet 0.970 0.960 0.980 0.980 0.980 0.940
ChemBERTa 0.757 0.540 0.529 0.987 0.661 0.986
TITAN 0.502 0.620 0.523 0.590 0.608 0.970
ERGO II 0.613 0.665 0.618 0.873 0.702 0.698
NetTCR 0.605 0.683 0.592 0.793 0.721 0.688
DlpTcr 0.520 0.716 0.307 0.017 0.030 0.514
Imrex 0.685 0.448 0.433 0.985 0.593 0.986
Table 4. Result of LANTERN and baseline models on Gen RN
Model AUC-ROC ACC RECALL PRECISION F1-SCORE AUC-PR
LANTERN 0.662 0.620 0.551 0.460 0.503 0.537
Smiles-Bert 0.556 0.513 0.498 0.443 0.468 0.491
TEINet 0.520 0.370 0.980 0.350 0.510 0.470
ChemBERTa 0.649 0.614 0.542 0.446 0.488 0.511
TITAN 0.525 0.660 0.537 0.516 0.523 0.551
ERGO II 0.512 0.643 0.503 0.516 0.534 0.503
NetTCR 0.534 0.524 0.557 0.563 0.609 0.524
DlpTcr 0.564 0.688 0.951 0.692 0.801 0.838
Imrex 0.660 0.685 0.113 0.731 0.194 0.518
Table 5. Result of LANTERN and baseline models on Original NA
Model AUC-ROC ACC RECALL PRECISION F1-SCORE AUC-PR
LANTERN 0.901 0.964 0.997 0.966 0.981 0.995
Smiles-Bert 0.658 0.705 0.637 0.890 0.821 0.928
TEINet 0.972 0.951 0.991 0.963 0.981 1.000
ChemBERTa 0.731 0.972 0.994 0.977 0.986 0.993
TITAN 0.820 0.980 0.999 0.990 0.990 1.000
ERGO II 0.623 0.689 0.621 0.857 0.603 0.725
NetTCR 0.593 0.602 0.615 0.813 0.498 0.632
DlpTcr 0.666 0.977 0.336 0.553 0.389 0.474
Imrex 0.760 0.954 0.987 0.965 0.976 0.988
Table 6. Result of LANTERN and baseline models on Original RN
Model AUC-ROC ACC RECALL PRECISION F1-SCORE AUC-PR
LANTERN 0.570 0.620 0.260 0.400 0.320 0.400
Smiles-Bert 0.573 0.602 0.521 0.763 0.702 0.817
TEINet 0.531 0.417 0.880 0.351 0.504 0.353
ChemBERTa 0.530 0.601 0.213 0.339 0.226 0.352
TITAN 0.501 0.660 0.413 0.562 0.431 0.347
ERGO II 0.521 0.505 0.537 0.689 0.531 0.625
NetTCR 0.535 0.511 0.547 0.621 0.509 0.662
DlpTcr 0.501 0.658 0.991 0.660 0.792 0.829
Imrex 0.562 0.657 0.026 0.387 0.046 0.377
@@
Comparison of Generated vs. Original Datasets
Comparing results between Gen NA and Original NA, and Gen RN and Original RN reveal that models trained on generated datasets generally performed better than those trained on original datasets. For instance, LANTERN achieves an AUC-ROC score of 0.88 on Gen NA versus 0.901 on Original NA, and 0.66 on Gen RN versus 0.57 on Original RN. This suggests that the novel negative sample generation method improved model training by providing more realistic and challenging scenarios, closely mimicking the complexities of real-world data. The generation of negative samples in Gen NA and Gen RN datasets allowed for a more balanced and comprehensive training process. This method ensures that the model encounters a variety of binding and non-binding interactions, enhancing its ability to distinguish between them. By incorporating a diverse range of negative samples, the model becomes more adept at recognizing subtle differences in binding patterns, leading to improved prediction accuracy and robustness.
Generally Challengeable Random Control Setting
@@
Embedding Cluster of Pretraining Model
Figure 3: t-SNE visualization of the embeddings from LANTERN model. — a. With pretrain of 0.857 h score. b. Without pretrain of 0.607 h score. c. Show frequency of the pretrained embeddings. d. Show frequency of the no pretrained embeddings.
@@
Quantitative Evaluation of Embedding Performance
Embedding Cluster of Pretraining Model
@@
To further quantify the performance of the embeddings learned by the pretraining model, we calculated several clustering performance metrics: Silhouette Index (SI) and Calinski-Harabasz Index (CHI). Additionally, we evaluated unsupervised metrics, including the Hopkins Score (HS) and Cluster Tendency Score (CTS) ref43. For SI and CH, we employed the K-means clustering method on the embeddings with k = 2. We put the algorithm for these metrics in the supplementary section 1. The results are illustrated in the Figure 4. In Figure 4, we observe the following metrics:
@@
Few-Shot Learning Capabilities of LANTERN
Figure 5: Few-Shot Learning performance of LANTERN — a. Performance comparison of models over different numbers of seen pairs. b. Performance comparison of models over different seen ratios.
@@
Discussion and Conclusion
LANTERN advances TCR-pMHC binding prediction by combining protein and molecular language models with cross-attention-based fusion. Its use of ESM and MolFormer embeddings enables the model to capture both sequence-level and chemical features, leading to strong performance in zero-shot and few-shot settings—particularly when predicting unseen epitopes. A key strength of LANTERN lies in its ability to generalize from limited data. Across various negative sampling strategies, the model consistently outperforms baseline methods, including both traditional deep learning and recent LLM-based approaches. The Multi-Head Cross-Attention module enhances interpretability and improves alignment between TCR and peptide features. While LANTERN performs well in both reference and random control settings, there is still room for improvement in distinguishing hard negatives. Future work may explore improved negative sampling, structural priors, or multimodal integration to further boost performance.
@@
Acknowledgments
So long and thanks for all the fish.
Evaluation Metrics
Hopkins Score
The Hopkins Score (HS) is a statistical test that measures the clustering tendency of a dataset, helping to determine whether the data points are more likely to form clusters or are uniformly distributed. The method begins by selecting a random subset of data points from the dataset. For each of these data points, a corresponding random point is generated within the same range as the data points. The next step involves calculating the distances between the nearest neighbors: $D_{T}$ represents the sum of the distances from each selected data point to its nearest neighbor within the subset, while $D_{R}$ represents the sum of the distances from each random point to its nearest data point. The Hopkins Score is then calculated using the formula $HS=\frac{D_{R}}{D_{R}+D_{T}}$. A Hopkins Score close to 1 indicates that the data has a strong clustering tendency, whereas a score around 0.5 suggests that the data points are uniformly distributed, showing no significant clustering.
Cluster Tendency Score
The Cluster Tendency Score (CTS) ref43 assesses how well region embeddings form clusters, under the assumption that clustered embeddings are more useful than dispersed ones. Inspired by the Hopkins test, the process starts with a set of $N_{S}=\{q_{i}\mid i=1,...,N_{S}\}$ embeddings. From $N_{S}$, a set of $N_{T}$ embeddings is selected as test points, denoted as $T=\{p_{i}\mid i=1,...,N_{T}\}$. For each test point $p_{i}$, its nearest neighbor $\hat{q}_{i}$ is found within $N_{S}\setminus\{q_{i}\}$. The sum of squared Euclidean distances between each test point $p_{i}$ and its nearest neighbor is calculated as $D_{T}=\sum_{i}=1^{N_{T}}d(p_{i},\hat{q}_{i}).$ Next, the same number of $N_{T}$ random points $U=\{u_{i}\mid i=1,...,N_{T}\}$ are sampled uniformly between the minimum and maximum values across each dimension of the embeddings in $N_{S}$. The sum of distances between each random point $u_{i}$ and its nearest neighbor $\hat{q}_{i}$ is calculated as $D_{R}=\sum_{i}=1^{N_{T}}d(u_{i},\hat{q}_{i}).$ The CTS is then defined as
$\text{CTS}=2\cdot\max\left(\frac{D_{R}}{D_{R}+D_{T}}-0.5,0\right),$
which normalizes the ratio between $D_{T}$ and $D_{R}$, with values ranging from 0 to 1. A higher CTS indicates stronger clustering, while if the embeddings are uniformly distributed, $D_{R}\approx D_{T}$, making $\frac{D_{R}}{D_{R}+D_{T}}\approx 0.5$, and CTS approaches 0.
Silhouette Index
The Silhouette Index (SI) is a metric that measures the quality of clustering by evaluating how similar each data point is to its own cluster compared to other clusters. For each data point $i$, the average distance $a(i)$ is computed between $i$ and all other points within the same cluster, which measures the cohesion or how tightly the point is clustered with its own group. Simultaneously, the average distance $b(i)$ is calculated between $i$ and all points in the nearest cluster that $i$ does not belong to, representing the separation or how distinct the cluster is from others. The Silhouette value for each data point is then calculated using the formula
$s(i)=\frac{b(i)-a(i)}{\max(a(i),b(i))},$
where the Silhouette score ranges from 1-1 to $1$.
A score close to $1$ indicates that the point is well-clustered, a score around 0 suggests that the point lies near the boundary between two clusters, and a negative score implies that the point might be misclassified, being closer to a different cluster than its own. The overall Silhouette Index is the mean of the silhouette values across all data points, providing an aggregate measure of clustering quality.
Calinski–Harabasz Index
The Calinski–Harabasz Index (CHI), also known as the Variance Ratio Criterion, is a metric used to assess the quality of clustering by evaluating the ratio of between-cluster dispersion to within-cluster dispersion.
The between-cluster dispersion $B$ measures how far apart the cluster centroids are from the overall centroid of the data and is calculated as
$B=\sum_{k=1}^{K}n_{k}\cdot\|c_{k}-c\|^{2},$
where $n_{k}$ is the number of points in cluster $k$, $c_{k}$ is the centroid of cluster $k$, and $c$ is the centroid of the entire dataset.
The within-cluster dispersion $W$ assesses how close the data points are to their respective cluster centroids and is calculated as
$W=\sum_{k=1}^{K}\sum_{i\in C_{k}}\|x_{i}-c_{k}\|^{2},$
where $x_{i}$ is a data point in cluster $k$, and $C_{k}$ is the set of points in cluster $k$.
The Calinski–Harabasz Index is then computed as
$CHI=\frac{B/(K-1)}{W/(N-K)},$
where $K$ is the number of clusters and $N$ is the total number of data points.
A higher CHI value indicates that the clusters are well-separated and internally compact, signifying a better clustering structure.